OpenStudy (anonymous):

Find the equation of the line that has the point (-4,1) and is parallel to y=3/4x

OpenStudy (anonymous):

y-1=3/4(x+4)

OpenStudy (anonymous):

lines are parallel,so slopes are equal

OpenStudy (anonymous):

lines are parallel so the whole equation will be same except the intercept. we know that equation of line is y = m.x + c here c is the intercept so the equation of the line parallel to the given line will be y = (3/4)x +c now as the line passes through (-4, 1) it will satisfy the equation so 1 = (3/4) -4 + c so 1 = -3 + c or c = 4 so the required equation is y = (3/4) x +4